OpenStudy (anonymous):

Name the function rule from the table given _______ X | Y ___|____ -4 | 4 -2 | 2 0 | 0 Option 1) Y= -3 Option 2) Y= -4 Option 3) Y= 2x Option 4) y=x X | Y ___|____ -9 | -18 0 | -9 9 | 0 Option 1) y= -2x Option 2) Y= X-9 Option 3) Y= X+9 Option 4) Y= 2x

OpenStudy (lexluth0r):

I'm not sure the answer is up there for the first one... :|

OpenStudy (lexluth0r):

and the second one is option 2

OpenStudy (lexluth0r):

but the first one I suspect should be y=x^2

OpenStudy (lexluth0r):

(x squared)

OpenStudy (freckles):

you mean y=-x ? y=x^2 wouldn't make since example f(x)=x^2 if x=-4 then f(-4)=16 which is not 4

OpenStudy (freckles):

sense*

OpenStudy (freckles):

@Iamnotsmart can you figure out the second problem? if you are made at making the rules, then just use the choices and see what points satisfy the choices

OpenStudy (freckles):

For example, x|y --- 2| 5 3| 6 4| 7 say we wanted to see if the rule for these points was f(x)=x+3 well f(2)=2+3=5 that is a point on the list and f(3)=3+3=6 and that is a point on the list and f(4)=4+3=7 which is the last point on the list f(x)=x+3 is definitely the rule for the list of points in my example but say you wanted to see if g(x)=x was a rule for the points I gave well g(2)=2 but we have that x=2 should correspond to y=5 so g(x)=x is definitely not the rule for the list of points in my example